OpenStudy (anonymous):

List all the subsets of {3, 4, 7}

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

I got it. Lol. Thanks though.

OpenStudy (skullpatrol):

Good job :)

OpenStudy (anonymous):

Thanks(:

OpenStudy (anonymous):

sub sets of {a,b,c} are \[\phi ,\left\{ a \right\},\left\{ b \right\},\left\{ c \right\},\left\{ a,b \right\},\left\{ b,c \right\} ,\left\{ c,a \right\} ,\left\{ a,b,c \right\} \]